You wish to buy enough honey to fill a 4.49 quart honey jar. At the grocery story you find that honey costs $3.00 per kilogram and it has a density of 1.50 grams per milliliter. How much will it cost in $ to buy enough honey to fill your 4.49 quart jar? (1 qt = 946 mL)
1
Expert's answer
2019-09-25T05:43:34-0400

First we should calculate the volume of honey required:


"V(honey, mL) = 4.49qt*946\frac{mL}{qt}=4248mL"


Now, using density we should found a mass of the honey required:


"M(honey, g)=4248mL*1.50\frac{g}{mL}=6372g(6.372kg)"


And now, it is the right time to found the price for the honey:


"P(honey, dollar)=6.372kg*3.00\frac{dollar}{kg}=19.12dollars"


Thus, we need 19 dollars and 12 cents to fill the jar with the honey.